OpenStudy (anonymous):

classify the system of equations and identify the number of solutions 2x+y=7 y+5= -2x

OpenStudy (dan815):

if you move the -2x over and the 5 over

OpenStudy (anonymous):

Is it consistent dependent infinte consistent independent one inconsistent none inconsistent dependent one

OpenStudy (dan815):

on the bottom equation, by adding 2x to both sides and subtracting 5 to both sides

OpenStudy (anonymous):

mhmm

OpenStudy (dan815):

what happens?

OpenStudy (anonymous):

idkkk

OpenStudy (dan815):

why did u say mhmm then

OpenStudy (anonymous):

haha idk

OpenStudy (anonymous):

easy .... kkkk

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

exchange value variable .

OpenStudy (dan815):

do you know about equations of line

OpenStudy (anonymous):

idont have enough time for this whats the answer

OpenStudy (dan815):

dont cheat

OpenStudy (anonymous):

oh really now

OpenStudy (anonymous):

study . kk 2x + y =7 -2x + y= 5

OpenStudy (anonymous):

now you can answer

OpenStudy (anonymous):

answer! haha ":(

OpenStudy (anonymous):

hahaha...

OpenStudy (anonymous):

:0

OpenStudy (dan815):

@ikram002p

OpenStudy (dan815):

@Squirrels @abb0t @Kainui

OpenStudy (ikram002p):

now as you can see 2x+y=7 2x+y=-5 are parallel ( they have same slope ) thus no solution . how ever if you wanna direct answer feed youre question to wolfram , it well do all the work :P http://www.wolframalpha.com/input/?i=2x%2By%3D7%2C+y%2B5%3D+-2x

OpenStudy (anonymous):

mhmm

OpenStudy (ikram002p):

-.- now what

OpenStudy (anonymous):

haha excuse mehh

OpenStudy (ikram002p):

study this girl , irs not that hard instead of saying mhmm and make ppl think that u dnt have brain ( sure u have brain ) but no one will help you unless u help urself .

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

uses the mind

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

feather

OpenStudy (bradely):

5 min i will give answer

OpenStudy (anonymous):

the hell u talking about you late

OpenStudy (ashleyisakitty):

Hi @Fani1996. Please come study with us again when you have decided to use Openstudy as a learning platform, not for a cheap answer. http://openstudy.com/code-of-conduct

OpenStudy (squirrels):

@Fani1996 noose yourslef